OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (7, -2) and has a slope of -3.

OpenStudy (anonymous):

y + 7 = -3(x - 2) y - 7 = -3(x + 2) y - 2 = -3(x + 7) y + 2 = -3(x - 7)

OpenStudy (anonymous):

use point-slope form: y - Y = m(x - X) where (X,Y) is a given point on the line and m is the slope since you're given (X,Y) = (7,-2) and m= -3 y - -2 = -3(x - 7) y + 2 = -3(x - 7) this is equal to option D.

OpenStudy (katrinakaif):

Plug it all into the equation: y-y1 = m (x-x1) with m = slope (-3) and your given points, (7,-2) equated with (x1,y1)